# Entropy-Based Decoding for Retrieval-Augmented Large Language Models

### v1

@@ Appendix A Appendix
A.1 Relations between Entropy and Performance
Figure 6 shows the average entropy values of tokens generated on the WebQ dataset by Llama2-7B and Llama2-13B when using different layers for contrast. For Llama2-7B, the entropy initially shows a decreasing trend and later an increasing trend among the tested layers, which correlates somewhat with the overall performance trend. For Llama2-13B, it was found that both the highest performance layer and the layer with the highest entropy are the last layer. This might explain why the performance of our entropy maximization-based layer selection is overall consistent with the performance of the last layer (see Figure 4).
A.2 Illustration of the False Positive Case

### v2

@@ Appendix A Appendix
A.1 Additional Implementation Details
As shown in Table 4, for both the dynamic JSD-based selection strategy and the proposed entropy-based strategy, we set the $\beta$ parameter to 5.0 for llama-2-7B and $\beta$ to 0.25 for other LLMs. This configuration ensures that contrastive decoding achieves superior performance across different LLMs. For the last-layer strategy, we found that setting a higher beta significantly severely reduces the performance of LLMs. Therefore, for this strategy, we set $\beta$ to 0.25 for all LLMs.
Method $\beta$
Naive w/ JSD 5.0 / 0.25
Naive w/ Last_Layer 0.25
Naive w/ Entropy 5.0/0.25
LeEns w/ JSD 5.0/0.25
LeEns w/ Last_Layer 0.25
LeEns w/ Entropy 5.0/0.25
Table 4: $\beta$ for different layer selection strategies.
A.2 Relations between Entropy and Performance
Figure 5 shows the average entropy values of tokens generated on the WebQ dataset by Llama2-7B and Llama2-13B when using different layers for contrast. For Llama2-7B, the entropy initially shows a decreasing trend and later an increasing trend among the tested layers, which correlates somewhat with the overall performance trend. For Llama2-13B, it was found that both the highest performance layer and the layer with the highest entropy are the last layer. This might explain why the performance of our entropy maximization-based layer selection is overall consistent with the performance of the last layer.
A.3 Illustration of the False Positive Case
@@ A.4 Prompt Format
A.5 Decoding Latency of Llama-2-13B
# of Docs 5 10 20
Naive 35.51 ($\times$1.00) 37.81 ($\times$1.00) 49.89 ($\times$1.00)
LeEns 39.14 ($\times$1.10) 43.54 ($\times$1.15) 58.84 ($\times$1.18)
CLeHe 40.24 ($\times$1.13) 45.28 ($\times$1.20) 60.79 ($\times$1.22)
Table 6: Decoding latency (ms/token) of Llama-2-13B based on the number of retrieved documents as context.
As shown in Table 6, when our methods are applied to the 13B model, the increase in decoding time is capped at factors of 1.22. This increase is considered acceptable given the superior performance our methods deliver.